OpenStudy (anonymous):

Write a word problem describing the time it takes to complete this activity individually and together. For example, if John takes 2 hours to mow his lawn and it takes his sister, Maria, 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together?

OpenStudy (anonymous):

If you call the two different times for completing the task x and y a general equation for solving all times can be written like this: \[x * y/(x + y)\] For the lawn mowing problem this gives us : \[2*4/(2+4) = 8/6 = 1.333... = 1 hour 20 minutes\]

OpenStudy (anonymous):

Thank you for letting me help you! Signed: \[Mangorox\]

OpenStudy (anonymous):

wait lol